# DE-BERT: Distance-Enhanced Early Exiting for BERT based on Prototypical Networks

### v1

@@ Abstract
Early exiting has demonstrated its effectiveness in accelerating the inference of pre-trained language models like BERT by dynamically adjusting the number of layers executed. However, most existing early exiting methods only consider local information from an individual test sample to determine their exiting indicators, failing to leverage the global information offered by sample population. This leads to suboptimal estimation of prediction correctness, resulting in erroneous exiting decisions. To bridge the gap, we explore the necessity of effectively combining both local and global information to ensure reliable early exiting during inference. Purposefully, we leverage prototypical networks to learn class prototypes and devise a distance metric between samples and class prototypes. This enables us to utilize global information for estimating the correctness of early predictions. On this basis, we propose a novel Distance-Enhanced Early Exiting framework for BERT (DE-BERT). DE-BERT implements a hybrid exiting strategy that supplements classic entropy-based local information with distance-based global information to enhance the estimation of prediction correctness for more reliable early exiting decisions. Extensive experiments on the GLUE benchmark demonstrate that DE-BERT consistently outperforms state-of-the-art models under different speed-up ratios with minimal storage or computational overhead, yielding a better trade-off between model performance and inference efficiency. Additionally, an in-depth analysis further validates the generality and interpretability of our method.
Index Terms:
Adaptive inference, early exiting, pre-trained language model (PLM), bidirectional encoder representations from transformers (BERT), prototypical network.I Introduction
@@
In recent years, large-scale pre-trained language models (PLMs), such as GPT [1], BERT [2], RoBERTa [3], and ALBERT [4], have brought significant improvements to natural language processing tasks. Despite the tremendous success, these transformer-based models still suffer from high memory and computational overhead in both training and inference. In particular, the prolonged inference latency hinders their deployment in edge devices or real-time scenarios. Besides, overthinking [5] poses another challenging issue for these large PLMs. It refers to the phenomenon where the model performs well in the intermediate layers for easy samples but experiences a decline in performance in the deeper layers, leading to redundant computation and performance degradation.
@@
To speed up the inference of PLMs, numerous attempts have been made by employing model compression to condense their sizes. These approaches encompass network pruning [6, 7, 8, 9, 10], weight quantization [11, 12, 13], knowledge distillation [14, 15, 16, 17], and weight sharing [18, 19]. These methods can effectively save memory usage and computational costs to accelerate inference. However, they often necessitate additional training costs to meet different acceleration requirements, leading to inflexible adjustments to the speed-up ratio.
@@
To address these issues, early exiting [20, 21, 22, 23, 24, 25, 26, 27] has been applied to accelerate the inference of PLMs. Unlike model compression, early exiting can easily adapt to different acceleration requirements by simply adjusting the threshold, without incurring additional training costs. In Fig. 1, we illustrate early exiting based on BERT, where an internal classifier is added to each intermediate layer of BERT. This enables the early exiting of samples when the early predictions from these internal classifiers are deemed sufficiently correct, eliminating the need to traverse the entire model. This strategy employs adaptive inference to deal with easy samples with shallow layers of BERT and process difficult samples with deeper layers. This approach effectively mitigates the overthinking problem and accelerates model inference while maintaining high accuracy.
@@
A typical implementation for early exiting is to devise an early exiting indicator that reflects the correctness of early predictions (i.e., prediction correctness) to establish the exiting criteria. According to the types of exiting indicators, there are mainly three early exiting strategies for the dynamic exiting of BERT models. The first strategy is confidence-based early exiting (e.g., Deebert [20], RightTool [28] and CascadeBERT [25]). It utilizes the model confidence, i.e., entropy or label score of the prediction probability distribution, to estimate the correctness of early predictions for exit decision-making. The second strategy is patience-based early exiting (e.g., PABEE [21] and LeeBERT [29]), relying on the cross-layer consistency to determine when to exit. Early exiting in the third type learns to generate an early exiting indicator that scores the prediction correctness (e.g., BERxiT [22]). However, a reliable exiting decision is to exit when the early predictions are sufficiently correct. Unfortunately, the absence of ground-truth labels during inference poses extreme challenges to estimating prediction correctness, which compromises the reliability of exiting decisions.
@@
In Fig. 1, we illustrate the confidence-based exit decision-making process for a sample in the QQP task. Notably, we observe that a high-confidence prediction made by a classifier is not necessarily correct [30]. This indicates that model confidence may not always accurately reflect prediction correctness, potentially leading to erroneous exiting decisions. In Fig. 2, we further compare the existing exiting strategies with the original backbone model, as well as the oracle which is an ideal model that can always enable each sample to exit at the shallowest internal classifier that provides a correct label prediction [28]. Essentially, the oracle serves as the upper bound for early exiting strategies as it provides the most accurate estimation of prediction correctness to achieve an optimal trade-off between model performance and efficiency. We can observe that the oracle outperforms the backbone model and existing exiting strategies by a large margin in both model performance and inference speed, which indicates significant room for improving the estimation of prediction correctness. We believe that the global information offered by the sample population can reflect the latent class information of test samples, which can be utilized to estimate the correctness of early predictions. However, the aforementioned strategies primarily focus on local information (e.g., entropy, consistency, and label scores) from an individual test sample to formulate their early exiting indicators, while neglecting global information provided by the sample population. Hence, insufficient information exacerbates the difficulty of estimating prediction correctness, resulting in unreliable exiting decisions.
@@
To bridge the gap, we explore the necessity of integrating both the local information from an individual test sample and the global information provided by the sample population to accurately estimate prediction correctness and formulate reliable early exiting during inference. Inspired by prototype learning [31] that involves learning class prototypes from a set of samples, we introduce a prototypical network paired with each internal classifier to learn class prototypes during training, as shown in Fig. 3LABEL:sub@fig:overview_ours. Then we devise a distance metric between samples and class prototypes to incorporate global information for estimating the correctness of early predictions. Accordingly, we propose a novel Distance-Enhanced Early Exiting framework for BERT models (DE-BERT) based on prototypical networks. DE-BERT leverages distance-based global information to enhance the classic entropy-based early exiting through information fusion. Concretely, in DE-BERT, each prototypical network is implemented with a simple linear layer and is jointly trained with other components under an elaborate distance-aware regularization (DAR). During the inference stage, we calculate distance scores between samples and class prototypes as a form of global information, in conjunction with the classic entropy-based local information, to facilitate estimating prediction correctness for reliable early exiting. To achieve this, a harmonic indicator integrating the distance-based global information and the entropy-based local information is proposed to formulate a hybrid exiting strategy. Our main contributions can be summarized as follows.
@@
We disclose that the performance limitation of existing early exiting methods primarily stems from the discrepancy between the exiting indicators and prediction correctness. We claim that this discrepancy arises from solely depending on local information from an individual test sample to make exiting decisions while neglecting the global information provided by the sample population.
@@
We propose a novel Distance-Enhanced Early Exiting framework for BERT (DE-BERT) with a hybrid exiting strategy incorporating both the traditional entropy-based local information and the distance-based global information to facilitate the estimation of prediction correctness for making reliable exiting decisions.
Extensive experiments on the GLUE benchmark demonstrate that our DE-BERT consistently outperforms state-of-the-art baselines under various speed-up ratios with minimal storage or computational overhead, yielding a superior trade-off between model performance and inference efficiency. Further analysis provides in-depth insights into the parameter sensitivity, the model’s interpretability and generality, storage and computational costs, as well as the contribution of each component.
The rest of this article is organized as follows. Section II provides a review of relevant work. Section III details our proposed DE-BERT framework. Extensive experiments and in-depth analysis are given in Section IV and Section V, respectively. Section VI discusses the experimental results and limitations of our framework, as well as directions for future work. Section VII concludes this article.
@@ II-A Static Model Compression
Model compression effectively compresses large PLMs to speed up inference while the compressed models remain static for all samples during inference. Specifically, network pruning [6, 7, 8, 9, 10] involves removing unnecessary weights and connections from the network. Weight quantization [11, 12, 13] reduces the precision of weight representations to save the computational and storage overhead of the model. Knowledge distillation [14, 15, 16, 17] transfers knowledge from a larger teacher model to a smaller student model. Weight sharing [18, 19] achieves model compression by sharing weights across different network components. However, compared with early exiting, these methods are less flexible to meet varying acceleration conditions, requiring additional training costs for speed-up ratio adjustments.
@@ II-B Dynamic Early Exiting
Early exiting, a parallel line of research for accelerating the inference of PLMs, is to dynamically stop inference for various input samples. DeeBERT [20], FastBERT [32], and RightTool [28] are pioneers in applying early exiting to accelerate the inference of BERT-style models. These studies employ the confidence level, such as entropy or label score of the prediction probability distribution, to estimate the correctness of early predictions for exit decision-making. On this basis, Lin et al. [33] extended confidence-based early exiting from classification tasks to sequence labeling tasks. To further improve the acceleration performance of early exiting models, one approach is to explore new exiting strategies for the inference stage. PABEE [21] uses the cross-layer consistency to determine early exiting. PCEE-BERT [23] employs a hybrid exiting strategy that combines confidence-based and patience-based strategies. BERxiT [22] learns to generate an exiting indicator that scores the correctness of early predictions. HASHEE [24] utilizes hash functions to assign each token to a fixed exiting layer, providing a novel static paradigm for early exiting. In addition, an alternative approach to improve acceleration performance is to strengthen the capability of internal classifiers through various techniques, such as model calibration, architecture refinement, and optimized training schemes. CascadeBERT [25] employs early exiting within multiple cascaded complete models, thus strengthening the representation capability of internal classifiers. Moreover, it formulates a difficulty-aware objective to calibrate the prediction probability distribution. LeeBERT [29] introduces a cross-layer distillation loss with learned weights to promote the training of internal classifiers. GPFEE [26] integrates both past and future states as inputs for each internal classifier, facilitating high-quality early predictions.
@@ II-C Prototypical Networks
Prototypical networks are first proposed in [31], which enable the learning of class prototypes from a set of training samples to classify the test sample based on the distances between the test sample and these prototypes. Prototypical networks provide an effective framework to learn a metric space for classification, which have been widely applied in few-shot image classification [34, 35] and image recognition tasks [36]. In recent years, prototypical networks have been further extended to the field of natural language processing, demonstrating successful applications in few-shot relation classification [37, 38] and named entity recognition tasks [39]. Additionally, prototypical networks have proven to be effective in the field of speech processing, particularly for speech recognition tasks in low-resource scenarios [40]. Besides the few-shot learning domain, prototype networks have also been successfully applied to domain adaptation [41] and anomaly detection [42].
Despite significant efforts dedicated to early exiting methods, there is still ample potential for improvement. The existing early exiting approaches rely solely on local information from an individual test sample to determine when to exit, overlooking the valuable global information conveyed by the sample population. This limited perspective can impact the accuracy of prediction correctness estimation, leading to unreliable exiting decisions. In contrast, inspired by prototype learning, we propose to leverage prototypical networks to enhance the classic entropy-based early exiting by incorporating global information based on a distance metric between the test sample and class prototypes. Our framework comprehensively considers both entropy-based local information and distance-based global information to improve the estimation of prediction correctness, thereby achieving a better trade-off between model performance and efficiency.
III The DE-BERT Framework
@@ III-A1 Problem Definition
Given a BERT-style model with layers, we denote the hidden states at the th layer as . To enable early exiting for the model’s inference on a classification task with classes, each of its intermediate layers is attached with an internal classifier to give an early prediction , i.e., a probability distribution over the classes.
@@ III-A2 Framework Overview
We propose a novel Distance Enhanced Early Exiting framework for BERT models (DE-BERT) based on the prototypical networks, which leverages the distance-based global information to enhance the classic entropy-based early exiting, aiming for reliable exiting decisions with sufficiently correct early predictions. Fig. 3LABEL:sub@fig:overview_ours provides an overview of our framework. Structurally, we first add one prototypical network (paired with one internal classifier) into each intermediate layer of BERT to learn class prototype representations. Each prototypical network consists of a simple linear layer that maps the hidden states of each layer to a metric space for distance calculation. We further propose distance-aware regularization (DAR) as the loss function to jointly train the prototypical networks with the BERT. Then, we devise a normalized distance ratio to provide global information for estimating the prediction correctness by calculating the distances between the sample representation and class prototypes. Finally, we propose a hybrid exiting strategy with a harmonic indicator that integrates both local information (entropy) and global information (distance ratio) to improve the estimation of prediction correctness.

### v2

# DE3-BERT: Distance-Enhanced Early Exiting for BERT Based on Prototypical Networks

@@ Abstract
Early exiting has demonstrated its effectiveness in accelerating the inference of pre-trained language models like BERT by dynamically adjusting the number of layers executed. However, most existing early exiting methods only consider local information from individual test samples to determine their exiting indicators, failing to leverage the global information offered by sample population. This leads to a suboptimal estimation of prediction correctness, yielding erroneous exiting decisions. To bridge the gap, we explore the necessity of effectively combining local and global information to ensure reliable early exiting during inference. Purposefully, we leverage prototypical networks to learn class prototypes and devise a distance metric between samples and class prototypes. This enables us to utilize global information for estimating the correctness of early predictions. On this basis, we propose a novel Distance-Enhanced Early Exiting framework for BERT (DE3-BERT). DE3-BERT implements a hybrid exiting strategy that supplements classic entropy-based local information with distance-based global information to enhance the estimation of prediction correctness for more reliable exiting decisions. Extensive experiments on the GLUE benchmark demonstrate that DE3-BERT consistently outperforms state-of-the-art models under different speed-up ratios with minimal storage or computational overhead, yielding a better trade-off between model performance and inference efficiency. Further analysis confirms the effectiveness of each component in our framework, as well as its generalization capability.
I Introduction
@@
In recent years, large-scale pre-trained language models (PLMs), such as GPT [radford2019language], BERT [bert, Shadi2024FakeND], RoBERTa [abs-1907-11692], and ALBERT [LanCGGSS20], have brought significant improvements to natural language processing tasks. Despite the tremendous success, these transformer-based models still suffer from high memory and computational overhead in both training and inference. In particular, the prolonged inference latency hinders their deployment in edge devices or real-time scenarios. Besides, overthinking [Overthinking] poses another challenging issue for these large PLMs. It refers to the phenomenon where the model performs well in intermediate layers for easy samples but experiences a decline in performance in deeper layers, leading to redundant computation and performance degradation.
@@
To speed up the inference of PLMs, numerous attempts have been made by employing model compression to condense their sizes. These approaches encompass network pruning [pruning4, pruning5], weight quantization [quantification2, quantification3], knowledge distillation [KD2, Yu2025ZeroCL, Ye2025CCDE], and weight sharing [weightsharing1, weightsharing2]. These methods can effectively save memory usage and computational costs to accelerate inference. However, they often necessitate additional training costs to meet different acceleration requirements, leading to inflexible adjustments to the speed-up ratio.
@@
To address these issues, early exiting [Deebert, Pabee, Berxit, PCEE, hashbased, NSP, twostageEE, li2021cascadebert, globalpast, ConsistentEE, COSEE] has been applied to accelerate the inference of PLMs. Unlike model compression, early exiting can easily adapt to different acceleration requirements by simply adjusting the threshold, without incurring additional training costs. In Fig. 1, we illustrate early exiting based on BERT, where an internal classifier is added to each intermediate layer of BERT. This enables the early exiting of samples when the early predictions from these internal classifiers are deemed sufficiently correct, eliminating the need to traverse the entire model. This strategy employs adaptive inference to deal with easy samples with shallow layers of BERT and process difficult samples with deeper layers. This approach effectively mitigates the overthinking problem and accelerates model inference while maintaining high accuracy.
@@
A typical implementation for early exiting is to devise an early exiting indicator that reflects the correctness of early predictions (i.e., prediction correctness) to establish the exiting criteria. According to the types of exiting indicators, there are mainly three early exiting strategies for the dynamic exiting of BERT models. The first strategy is confidence-based early exiting (e.g., DeeBERT [Deebert], RightTool [Righttool], and CascadeBERT [li2021cascadebert]). It utilizes the model confidence, i.e., entropy or label score of the prediction probability distribution, to estimate the correctness of early predictions for exit decision-making. The second strategy is patience-based early exiting (e.g., PABEE [Pabee] and LeeBERT [zhu2021leebert]), relying on the cross-layer consistency to determine when to exit. Early exiting in the third type learns to generate an early exiting indicator for making exiting decisions (e.g., BERxiT [Berxit] and ConsistentEE [ConsistentEE]). However, a reliable exiting strategy is to exit the sample when its early prediction is sufficiently correct. Unfortunately, the absence of ground-truth labels during inference poses extreme challenges to estimating prediction correctness, which compromises the reliability of exiting decisions.
@@
In Fig. 1, we illustrate the confidence-based exit decision-making process for a sample in the QQP task. Notably, we observe that a high-confidence prediction made by a classifier is not necessarily correct [Z22ood]. This indicates that model confidence may not always accurately reflect prediction correctness, potentially leading to erroneous exiting decisions. In Fig. 2, we further compare the existing exiting strategies with the original backbone model, as well as the oracle which is an ideal model that can always enable each sample to exit at the shallowest internal classifier that provides a correct label prediction [Righttool]. Essentially, the oracle serves as the upper bound for early exiting strategies as it provides the most accurate estimation of prediction correctness to achieve an optimal trade-off between model performance and efficiency. We can observe that the oracle outperforms the backbone model and existing exiting strategies by a large margin in both model performance and inference speed, which indicates significant room for improving the estimation of prediction correctness. We believe that the global information offered by the sample population can reflect the latent class information of test samples, which can be utilized to estimate the correctness of early predictions. However, the aforementioned strategies primarily focus on local information (e.g., entropy, consistency, and label scores) from individual test samples to formulate their early exiting indicators, while neglecting global information provided by the sample population. Hence, insufficient information exacerbates the difficulty of estimating prediction correctness, resulting in unreliable exiting decisions.
@@
To bridge the gap, we explore the necessity of integrating both the local information from individual test samples and the global information provided by the sample population to accurately estimate prediction correctness and formulate reliable early exiting during inference. Inspired by prototype learning [snell2017prototypical] that involves learning class prototypes from a set of samples, we introduce a prototypical network paired with each internal classifier to learn class prototypes during training, as shown in Fig. 3LABEL:sub@fig:overview_ours. Then we devise a distance metric between samples and class prototypes to incorporate global information for estimating the correctness of early predictions. Accordingly, we propose a novel Distance-Enhanced Early Exiting framework for BERT (DE3-BERT) based on prototypical networks. DE3-BERT leverages distance-based global information to enhance the classic entropy-based early exiting through information fusion. Concretely, in DE3-BERT, each prototypical network is implemented with a simple linear layer and is jointly trained with other components under an elaborate distance-aware regularization (DAR). During inference, we calculate distance scores between samples and class prototypes, and then integrate the distance-based global information with the classic entropy-based local information to enhance prediction correctness estimation for more reliable early exiting. To achieve this, a harmonic indicator that integrates global and local information is proposed to formulate a hybrid exiting strategy. Our contributions can be summarized as follows.
@@
We disclose that the performance limitation of existing early exiting methods primarily stems from the discrepancy between exiting indicators and prediction correctness. We claim that this discrepancy arises from solely depending on local information from individual test samples to make exiting decisions while neglecting the global information provided by the sample population.
@@
We propose a novel Distance-Enhanced Early Exiting framework for BERT (DE3-BERT) with a hybrid exiting strategy considering both the traditional entropy-based local information and the distance-based global information to facilitate the estimation of prediction correctness for making reliable exiting decisions.
Extensive experiments on the GLUE benchmark demonstrate that our DE3-BERT achieves a superior performance-efficiency trade-off compared to state-of-the-art baseline methods, with minimal additional storage or computational overhead. Further analysis validates the effectiveness of each component within our framework and confirms the necessity of incorporating global information to ensure accurate estimation of prediction correctness. Additionally, our method shows insensitivity to parameter selection and demonstrates strong generality across out-of-distribution (OOD) data, different languages, and various backbone models.
@@ II-A Static Model Compression
Model compression effectively compresses large PLMs to speed up inference while the compressed models remain static for all samples during inference. Specifically, network pruning [pruning4, pruning5] involves removing unnecessary weights and connections from the network. Weight quantization [quantification2, quantification3] reduces the precision of weight representations to save the computational and storage overhead of the model. Knowledge distillation [KD2, Yu2025ZeroCL, Ye2025CCDE] transfers knowledge from a larger teacher model to a smaller student model. Weight sharing [weightsharing1, weightsharing2] achieves model compression by sharing weights across different network components. However, compared with early exiting, these methods are less flexible to meet varying acceleration conditions, requiring additional training costs for speed-up ratio adjustments.
@@ II-B Dynamic Early Exiting
Early exiting, a parallel line of research for accelerating the inference of PLMs, is to dynamically stop inference for various input samples. DeeBERT [Deebert], FastBERT [Fastbert], and RightTool [Righttool] are pioneers in applying early exiting to accelerate the inference of BERT-style models. These studies employ the confidence level, such as entropy or label score of the prediction probability distribution, to estimate the correctness of early predictions for exit decision-making. On this basis, Lin et al. [EEforSequenceLabeling] extended confidence-based early exiting from classification tasks to sequence labeling tasks. CeeBERT [ceebert] introduces an online learning algorithm to tackle threshold selection for confidence-based early exiting in cross-domain scenarios. To further improve the acceleration performance of early exiting models, one approach is to explore new exiting strategies for the inference stage. PABEE [Pabee] uses the cross-layer consistency to determine early exiting. PCEE-BERT [PCEE] employs a hybrid exiting strategy that combines confidence-based and patience-based strategies. BERxiT [Berxit] and ConsistentEE [ConsistentEE] leverage neural networks to generate exiting indicators for making exiting decisions. HASHEE [hashbased] utilizes hash functions to allocate each token to a designated exiting layer, providing a novel static paradigm for early exiting. In addition, an alternative approach to improve acceleration performance is to strengthen the capability of internal classifiers through various techniques, such as model calibration, architecture refinement, and optimized training schemes. CascadeBERT [li2021cascadebert] employs early exiting within multiple cascaded complete models, thus strengthening the representation capability of internal classifiers. Moreover, it incorporates a difficulty-aware objective to calibrate the prediction probability distribution. LeeBERT [zhu2021leebert] introduces a cross-layer distillation loss with learned weights to promote the training of internal classifiers. GPFEE [globalpast] integrates both past and future states as inputs for each internal classifier, facilitating high-quality early predictions. ConsistentEE [ConsistentEE] minimizes the classification loss for each sample solely at its exiting layer, thus enhancing consistency between training and testing to improve the capability of each classifier. COSEE [COSEE] assigns sample-wise loss weights for each classifier to bridge the gap between training and testing while enabling flexible adjustments to the speed-up ratio.
@@ II-C Prototypical Networks
Prototypical networks are first proposed in [snell2017prototypical], which enable the learning of class prototypes from a set of training samples to classify the test sample based on the distances between the test sample and these prototypes. Prototypical networks provide an effective framework to learn a metric space for classification, which have been widely applied in few-shot image classification [imageclassification1, imageclassification2] and image recognition tasks [imagerecognition]. In recent years, prototypical networks have been further extended to the field of natural language processing, demonstrating successful applications in few-shot relation classification [relationclassification1, relationclassification2] and named entity recognition tasks [ner]. Additionally, prototypical networks have proven to be effective in the field of speech processing, particularly for speech recognition tasks in low-resource scenarios [speechrecognition]. Besides the few-shot learning domain, prototypical networks have also been successfully applied to domain adaptation [PrototypeDomainAdaption] and anomaly detection [PrototypeAnomalyDetection].
Despite significant efforts dedicated to early exiting methods, there is still ample potential for improvement. The existing early exiting approaches rely solely on local information from individual test samples to determine when to exit, overlooking the valuable global information conveyed by the sample population. This limited perspective can impact the accuracy of prediction correctness estimation, leading to unreliable exiting decisions. In contrast, inspired by prototype learning, we propose to leverage prototypical networks to enhance the classic entropy-based early exiting by incorporating global information based on a distance metric between the test sample and class prototypes. Our framework comprehensively considers both entropy-based local information and distance-based global information to improve the estimation of prediction correctness, thereby achieving a better trade-off between model performance and efficiency.
III The DE3-BERT Framework
@@ III-A1 Problem Definition
Given a BERT-style model with layers, we denote the hidden states at the th layer as . To enable early exiting for the model’s inference on a classification task with classes, each of its intermediate layers is attached with an internal classifier where to give an early prediction , i.e., a probability distribution over the classes.
@@ III-A2 Framework Overview
We propose a novel Distance-Enhanced Early Exiting framework for BERT (DE3-BERT) based on the prototypical networks, which leverages the distance-based global information to enhance the classic entropy-based early exiting, aiming for reliable exiting decisions with sufficiently correct early predictions. Fig. 3LABEL:sub@fig:overview_ours provides an overview of our framework. Structurally, we first add one prototypical network (paired with one internal classifier) into each intermediate layer of BERT to learn class prototype representations. Each prototypical network consists of a simple linear layer that maps the hidden states of each layer to a metric space for distance calculation. We further propose distance-aware regularization (DAR) as the loss function to jointly train the prototypical networks with the BERT. Then, we devise a normalized distance ratio to provide global information for estimating the prediction correctness by calculating the distances between the sample representation and class prototypes. Finally, we propose a hybrid exiting strategy with a harmonic indicator that integrates both local information (entropy) and global information (distance ratio) to improve the estimation of prediction correctness.